OpenStudy (anonymous):

A rectangular tank has a length of 3 ft, a width of 4 ft, and a height of 5 ft. The tank is half filled with water. If a metal cube with the sides that are 1 ft long is submerged in the water, by how many INCHES will the level of the water rise?

OpenStudy (anonymous):

MMMM... In... Christian please XD

OpenStudy (anonymous):

60 = volume of tank

OpenStudy (dumbcow):

Volume of tank = 3*4*5 = 60 Tank is half filled so there is 30 cubic feet of water height of water is 2.5ft Volume of cube = 1^3 = 1 height of cube is 1 ft, so volume of water in first foot is displaced by 1. New volume for first foot of water = (3*4*1)-1 = 11 That leaves 19 cubic feet of water left 19 = 3*4*h h = 19/12 total height of water is 19/12 +1 = 31/12 Subtract 2.5 or 30/12 31/12 - 30/12 = 1/12 water rose 1/12 feet or 1 inch